OpenStudy (anonymous):

Verify the identity. cos 4x + cos 2x = 2 - 2 sin2 2x - 2 sin2 x

OpenStudy (anonymous):

Is your equation written as such: cos(4x)+cos(2x)=2-2sin(2x)^2-2sin(x)^2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Do you know the double angle identity for cos(2x)?

OpenStudy (shamim):

\[\cos 4x+\cos 2x=\cos2(2x)+1-2\sin ^{2}x=1-2\sin ^{2}2x+1-2\sin ^{2}x\]

OpenStudy (anonymous):

no

OpenStudy (shamim):

\[=2-2\sin ^{2}2x-2\sin ^{2}x\]

OpenStudy (anonymous):

Okay, so replace cos(2x) with 1-2sinx^2 and cos(4x) with 1-2sin(2x)^2

OpenStudy (shamim):

ya

OpenStudy (anonymous):

Then you have 1-2sin(x)^2+1-2sin(2x)^2=right side.

OpenStudy (anonymous):

simplify the variables and you have the left side= to right side.

OpenStudy (anonymous):

Shamim is that the answer?

OpenStudy (shamim):

i c

OpenStudy (shamim):

i m happy to solve this problem

OpenStudy (shamim):

ya its my answer

OpenStudy (shamim):

lhs=rhs